(NPV calculation​) Calculate the NPV given the following free cash​ flows,

YEAR   CASH FLOWS
0   -50,000
1   30,000
2   30,000
3   30,000
4   -30,000
5   30,000
6   30,000

if the appropriate required rate of return is 12 percent. Should the project be​ accepted? What is the​ project's NPV​?
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=30,000/1.12+30,000/1.12^2+30,000/1.12^3+30,000/1.12^5+30,000/1.12^6

=104276.68

Present value of outflows=50,000+30,000/1.12^4

=69065.54

NPV=Present value of inflows-Present value of outflows

=104276.68-69065.54

=$35211.14(Approx).

Hence since NPV is positive;project must be accepted.